OpenStudy (anonymous):

Which of the following equations represents a line perpendicular to 6x - 5y = 8? Question 3 options: -6x + 5y = 10 -5x + 6y = 8 6x + 5y = 1 5x + 6y = 12

OpenStudy (anonymous):

y = 6x/5-8/5 The perpendicular line is going to have a slope that is the "negative inverse" Therefore, m=-5x/6 Just by looking at your options, you have two avaliable options that could be possible answers the one with -5x, and the one with 5x. I'll check the last one b/c I think that's the possible answer. 5x+6y = 12 6y = -5x +12 y=-5x/6 +12/6 y=-5x/6 +2 The last option DOES have a slope of m=-5/6. Doesn't matter what the y-intercept is. This line is perpendicular to the given line, so this is your answer.